# HOG-Diff: Higher-Order Guided Diffusion for Graph Generation

### v2

@@ 3.3 Theoretical Analysis
$\mathbb{E}\left[\ell^{(k)}(\bm{\theta}_{i})\right]\leq\left(1-\frac{b\mu^{2}}{\beta N(\beta N^{2}+\mu(b-1))}\right)^{i}\ell^{(k)}\left(\bm{\theta}_{0}\right),$ (9)
@@
Under appropriate Lipschitz and boundedness assumptions, the reconstruction error of HOG-Diff satisfies the following bound:
$\mathcal{E}(0)\leq\alpha(0)\exp{\int_{0}^{\tau_{1}}\gamma(s)}\mathop{}\!\mathrm{d}{s},$ (10)
where $\alpha(0)=C^{2}\ell^{(1)}(\bm{\theta})\int_{0}^{\tau_{1}}g_{1,s}^{4}\mathop{}\!\mathrm{d}{s}+C\mathcal{E}(\tau_{1})\int_{0}^{\tau_{1}}h_{1,s}^{2}\mathop{}\!\mathrm{d}{s}$, $\gamma(s)=C^{2}g_{1,s}^{4}\|\bm{s}_{\bm{\theta}}(\cdot,s)\|_{\mathrm{lip}}^{2}+C\|h_{1,s}\|_{\mathrm{lip}}^{2}$, and $h_{1,s}=\theta_{s}\left(1+\frac{2}{e^{2\bar{\theta}_{s:\tau_{1}}}-1}\right)$. Furthermore, we can derive that the reconstruction error bound of HOG-Diff is sharper than that of classical graph generation models.
The theorems above rely primarily on mild assumptions, such as smoothness and boundedness, without imposing strict conditions like the target distribution being log-concave or satisfying the log-Sobolev inequality. Their formal statements and detailed proofs are postponed to App. A.
@@ 4 Experiments
We assess HOG-Diff against state-of-the-art baselines in both molecular and generic graph generation. Ablation studies are further conducted to analyze the impact of different topological guides. Complexity analysis and experimental settings are deferred to Apps. D and E, while App. F presents further results, including diffusion domain analysis, large-scale SBM experiments, the rationale for filtering choices, variance statistics, and visualizations.
Table 1: Comparison of different methods based on molecular datasets. We report the mean of 3 different runs. The best results for the first three metrics are highlighted in bold.
@@
Val. w/o
@@
[…]
Val. w/o
[…]
GraphAF 0.021 5.625 74.43 88.64 86.59 0.044 16.023 68.47 98.64 99.99
GraphDF 0.064 10.928 93.88 98.58 98.54 0.177 33.546 90.61 99.63 100.00
GraphArm 0.002 1.220 90.25 95.62 70.39 0.055 16.260 88.23 99.46 100.00
MiCaM 0.001 1.045 99.93 93.89 83.25 0.166 31.495 100.00 88.48 99.98
GraphEBM 0.030 6.143 8.22 97.90 97.01 0.212 35.471 5.29 98.79 100.00
SPECTRE 0.163 47.960 87.30 35.70 97.28 0.109 18.440 90.20 67.05 100.00
GSDM 0.003 2.650 99.90 - - 0.017 12.956 92.70 - -
EDP-GNN 0.005 2.680 47.52 99.25 86.58 0.049 16.737 82.97 99.79 100.00
GDSS 0.003 2.900 95.72 98.46 86.27 0.019 14.656 97.01 99.64 100.00
DiGress 0.0005 0.360 99.00 96.66 33.40 0.082 23.060 91.02 81.23 100.00
MoFlow 0.017 4.467 91.36 98.65 94.72 0.046 20.931 63.11 99.99 100.00
CatFlow - 0.441 99.81 99.95 - - 13.211 99.95 99.99 -
DeFoG 0.0005 0.268 99.26 96.61 72.57 0.002 2.030 94.97 99.98 100.00
HOG-Diff 0.0003 0.172 98.74 97.10 75.12 0.001 1.633 98.56 99.96 99.53
@@ 4.1 Molecule Generation
Molecular design is a prominent application of graph generation. We conduct evaluations on two well-known molecular datasets: QM9 (Ramakrishnan et al., 2014) and ZINC250k (Irwin et al., 2012). Intermediate higher-order skeletons are extracted using 2-cell complex filtering; the rationale for this choice is discussed in Sec. F.3. We evaluate the quality of 10,000 generated molecules with five standard metrics as in Jo et al. (2022): Neighborhood Subgraph Pairwise Distance Kernel (NSPDK) MMD (Costa & Grave, 2010), Fréchet ChemNet Distance (FCD) (Preuer et al., 2018), Validity without correction (Val. w/o corr.), Uniqueness (Uni.), and Novelty (Nov.) (Jo et al., 2022).
For benchmarking, we include various representative molecular generation models. Autoregressive models include GraphAF (Shi et al., 2020), GraphDF (Luo et al., 2021), and GraphArm (Kong et al., 2023); the fragment-based MiCaM (Geng et al., 2023) creatively leverage motif information. In contrast, the remaining methods adopt a one-shot generation paradigm: GraphEBM serves as an energy-based model, SPECTRE (Martinkus et al., 2022) and GSDM (Luo et al., 2023a) incorporate spectral conditioning within GAN and diffusion frameworks, respectively, while EDP-GNN (Niu et al., 2020), GDSS (Jo et al., 2022), and DiGress (Vignac et al., 2023) represent diffusion-based graph generation models. We also compare against advanced flow-based models, including MoFlow (Zang & Wang, 2020), CatFlow (Eijkelboom et al., 2024), and DeFoG (Qin et al., 2025).
We visualize the molecule generation process in Fig. 3 with more examples deferred to Sec. F.5. It can be observed that our model explicitly preserves higher-order structures during the generation process. Tab. 1 indicates that HOG-Diff consistently outperforms both auto-regressive and one-shot models. Notably, the dramatic decrease in NSPDK and FCD implies that HOG-Diff is able to generate molecules with data distributions close to those of real molecules in both chemical and graph spaces. Additional results on validity and variance are provided in Sec. F.4.
@@ 4.2 Generic Graph Generation
To display the topology distribution learning ability, we assess HOG-Diff over four common generic graph datasets: Community-small, Ego-small, Enzymes, and a larger-scale stochastic block model (SBM) dataset. The SBM dataset is presented separately in Sec. F.2 due to its distinct evaluation protocol. Intermediate higher-order skeletons are obtained via 3-simplicial complex filtering. We employ the same train/test split as Jo et al. (2022) for a fair comparison. Maximum mean discrepancy (MMD) is used to quantify the distribution differences across key graph statistics, including degree (Deg.), clustering coefficient (Clus.), and 4-node orbit counts (Orbit). A lower MMD signifies a closer alignment between the generated and evaluation datasets, suggesting superior generative performance. We also report the average MMD across all metrics as an overall indicator.
@@
We compare the following graph generative models: DeepGMG (Li et al., 2018) and GraphRNN (You et al., 2018) are autoregressive models, while GraphVAE (Simonovsky & Komodakis, 2018), GNF (Liu et al., 2019), and GPrinFlowNet (Mo et al., 2024) are one-shot models. GraphAF, GraphDF, EDP-GNN, SPECTRE, GDSS, and DiGress are previously explained. The results in Tab. 2 verify that HOG-Diff is not only suitable for molecular generation but also proficient in generic graph generation, demonstrating its ability to effectively capture the intricate topological interdependencies.
4.3 Ablations: Topological Guide Analysis
During the experiments, we observe that HOG-Diff exhibits superior performance on complex datasets such as QM9 and ZINC250k, but comparatively modest results on the Ego-small dataset. Statistics and visualizations in Apps. E and F.5 reveal that Ego contains the fewest higher-order structures among the datasets analyzed, suggesting that the choice of guide plays a pivotal role in the effectiveness of generation. To validate this hypothesis, we conduct further ablations using different types of topological information as guides.
@@ 4.3 Ablations: Topological Guide Analysis
Val. w/o
[…]
Noise 0.0015 0.829 91.52
QM9 Periph. 0.0009 0.305 97.58
Cell 0.0003 0.172 98.74
ZINC
250k
Noise 0.002 2.665 96.78
Periph. 0.002 2.641 97.93
Cell 0.001 1.633 98.56
In the sampling procedure, we further evaluate peripheral structures (Periph.), obtained by removing cell components, as guides. As shown in Tab. 3, both peripheral and noise guides perform worse than cell-based guides, providing empirical support for the tighter reconstruction error bound established in Theorem 4. These results indicate that certain topologies, particularly cells, are more effective in guiding generation, highlighting the importance of selecting appropriate topological structures to steer toward meaningful outputs. Moreover, this finding suggests that guides could serve as diagnostic tools for assessing whether specific topologies are essential to the architecture. Systematic analysis of different guides promises to enhance interpretability, clarify how topologies affect generation, and inform the design of more effective graph generative models.
5 Related Works
@@
Deep Generative Models
Graph generative models have made great progress by exploiting the capacity of deep neural networks. These models typically generate nodes and edges either in an autoregressive manner or simultaneously, utilizing techniques such as variational autoencoders (VAE) (Jin et al., 2018; Simonovsky & Komodakis, 2018), recurrent neural networks (RNN) (You et al., 2018), normalizing flows (Zang & Wang, 2020; Shi et al., 2020; Luo et al., 2021), and generative adversarial networks (GAN) (Martinkus et al., 2022).
@@
A breakthrough in graph generative models has been marked by the recent progress in diffusion-based generative models (Niu et al., 2020). Recent models employ various strategies to enhance the generation of complex graphs, including capturing node-edge dependency (Jo et al., 2022), addressing discretization challenges (Vignac et al., 2023; Huang et al., 2023), exploiting low-to-high frequency generation curriculum (Mo et al., 2024), and improving computational efficiency through low-rank diffusion processes (Luo et al., 2023a). CatFlow (Eijkelboom et al., 2024) and DeFoG (Qin et al., 2025) adopt flow matching as an alternative to diffusion, achieving more efficient generation. Recent studies have also enhanced diffusion-based generative models by incorporating diffusion bridge processes, i.e., processes conditioned on the endpoints (Wu et al., 2022; Boget et al., 2024; Jo et al., 2024). Despite these advances, existing methods either overlook or inadvertently disrupt higher-order structures during graph generation, or struggle to model the topological properties, as denoising the noisy samples does not explicitly preserve the intricate structural dependencies required for generating realistic graphs.
@@ 6 Conclusion
We introduce HOG-Diff, a coarse-to-fine generation framework that explicitly exploits higher-order graph topology. It decomposes the complicated graph generation process into easier-to-learn sub-steps, which are implemented using a GOU bridge process. Our theoretical analysis justifies the effectiveness of HOG-Diff over classical diffusion-based approaches, which is further validated by superior experimental results on both molecular and generic graph generation tasks. This work is a key step in topological diffusion models, highlighting the impact of higher-order features absent in data and opening ample room for future work.

### v3

@@ 3.3 Theoretical Analysis
$\mathbb{E}\left[\ell^{(k)}(\bm{\theta}_{i})\right]\leq\left(1-\frac{b\mu^{2}}{\beta N(\beta N^{2}+\mu(b-1))}\right)^{i}\ell^{(k)}\left(\bm{\theta}_{0}\right),$
@@
Under standard Lipschitz and boundedness assumptions, the reconstruction error of HOG-Diff satisfies the following bound:
$\mathcal{E}_{\mathrm{hog}}(0)\leq A_{\mathrm{hog}}\,e^{T_{\mathrm{hog}}},$ (10)
where $A_{\mathrm{hog}}=C^{2}\sum_{k=1}^{K}\ell^{(k)}(\bm{\theta})\int_{\tau_{k-1}}^{\tau_{k}}g_{k,s}^{4}\mathrm{d}s$ and $T_{\mathrm{hog}}=\sum_{k=1}^{K}\int_{\tau_{k-1}}^{\tau_{k}}\gamma_{k}(s)\mathrm{d}s+C\sum_{k=1}^{K-1}\int_{\tau_{k-1}}^{\tau_{k}}h_{k,s}^{2}\mathrm{d}s$. Moreover, under the same assumptions, HOG-Diff admits a tighter reconstruction error bound than classical diffusion models.
The theorems above rely primarily on mild assumptions, such as smoothness and boundedness, without imposing strict conditions like the target distribution being log-concave or satisfying the log-Sobolev inequality. Their formal statements and detailed proofs are postponed to App. A. We experimentally verify these Theorems in Sec. 4.4.
@@ 4 Experiments
Table 1: Comparison of different methods on molecular datasets. The best results are highlighted in bold. The full version with more baselines, metrics, and standard deviations is provided in Sec. F.1.
@@
Val. $\uparrow$ Uni. $\uparrow$ Nov. $\uparrow$ FCD $\downarrow$ NSPDK $\downarrow$ Val. $\uparrow$ Uni. $\uparrow$ Nov. $\uparrow$ FCD $\downarrow$ NSPDK $\downarrow$
GraphAF 74.43 88.64 86.59 5.625 0.021 68.47 98.64 99.99 16.023 0.044
GraphDF 93.88 98.58 98.54 10.928 0.064 90.61 99.63 100.00 33.546 0.177
GraphArm 90.25 95.62 70.39 1.220 0.002 88.23 99.46 100.00 16.260 0.055
MiCaM 99.93 93.89 83.25 1.045 0.001 100.00 88.48 99.98 31.495 0.166
SPECTRE 87.30 35.70 97.28 47.960 0.163 90.20 67.05 100.00 18.440 0.109
GSDM 99.90 - - 2.650 0.003 92.70 - - 12.956 0.017
EDP-GNN 47.52 99.25 86.58 2.680 0.005 82.97 99.79 100.00 16.737 0.049
GDSS 95.72 98.46 86.27 2.900 0.003 97.01 99.64 100.00 14.656 0.019
DiGress 99.00 96.66 33.40 0.360 0.0005 91.02 81.23 100.00 23.060 0.082
MoFlow 91.36 98.65 94.72 4.467 0.017 63.11 99.99 100.00 20.931 0.046
Cometh 99.59 96.75 72.06 0.248 0.0005 - - - - -
CatFlow 99.81 99.95 - 0.441 - 99.95 99.99 - 13.211 -
DeFoG 99.26 96.61 72.57 0.268 0.0005 94.97 99.98 100.00 2.030 0.002
HOG-Diff 98.74 97.10 75.12 0.172 0.0003 98.56 99.96 99.53 1.633 0.001
@@
MOSES DiGress 85.7 100.0 95.0 1.19 97.1
DisCo 88.3 100.0 97.7 1.44 95.6
Cometh 90.5 99.9 92.6 1.27 99.1
DeFoG 92.8 99.9 92.1 1.95 98.9
HOG-Diff 99.7 100.0 89.1 0.94 96.7
@@
GuacaMol DiGress 85.2 100.0 99.9 68.0 92.9
DisCo 86.6 86.6 86.5 59.7 92.6
Cometh 98.9 98.9 97.6 72.7 96.7
DeFoG 99.0 99.0 97.9 73.8 97.7
HOG-Diff 99.1 100.0 97.0 78.5 96.9
Figure 3: (left) Large molecule generation performance. Best results are highlighted in bold. Full versions are presented in Sec. F.2. (right) Generative trajectories progressing from left to right, illustrating that HOG-Diff preserves topological integrity across intermediate steps.
We assess HOG-Diff against state-of-the-art baselines in both molecular and generic graph generation. Ablation studies are further conducted to analyze the impact of different topological guides. Complexity analysis and experimental settings are deferred to Appendices D and E, while App. F reports additional results, including scalability to large benchmarks, variance statistics, diffusion-domain analysis, the rationale for filtering choices, and visualizations.
@@ 4.1 Molecule Generation
Molecular design is a prominent application of graph generation. We evaluate our method on four widely used molecular benchmarks: QM9 (Ramakrishnan et al., 2014), ZINC250k (Irwin et al., 2012), and the large-scale MOSES (Polykovskiy et al., 2020) and GuacaMol (Brown et al., 2019) benchmarks. In contrast to many prior works that report results on only a subset of these datasets, our evaluation provides comprehensive coverage. Intermediate higher-order skeletons are extracted using 2-cell complex filtering; the rationale for this choice is discussed in Sec. F.4. For QM9 and ZINC250k, we follow the dataset split and evaluation protocols of Jo et al. (2022). For MOSES and GuacaMol, we adopt their official splits and benchmark-specific metrics. We report Validity without correction (Val.), Uniqueness (Uni.), Novelty (Nov.), Fréchet ChemNet Distance (FCD) (Preuer et al., 2018), NSPDK (Costa and Grave, 2010), Filters, and KL divergence (Brown et al., 2019).
We benchmark against a diverse set of representative molecular generators. Autoregressive approaches include GraphAF (Shi et al., 2020), GraphDF (Luo et al., 2021), GraphArm (Kong et al., 2023), and the fragment-based MiCaM (Geng et al., 2023). The remaining methods adopt a one-shot paradigm: SPECTRE (Martinkus et al., 2022) incorporates spectral conditioning within a GAN framework, while GDSS (Jo et al., 2022), DiGress (Vignac et al., 2023), and Cometh (Siraudin et al., 2025) are diffusion-based models. We also compare against advanced flow-based models, namely MoFlow (Zang and Wang, 2020), CatFlow (Eijkelboom et al., 2024), and DeFoG (Qin et al., 2025).
Sampling Quality. Tab. 1 and Fig. 3(left) show that HOG-Diff generally outperforms both auto-regressive and one-shot models. Notably, the dramatic decrease in NSPDK and FCD implies that HOG-Diff is able to generate molecules with data distributions close to those of real molecules in both graph and chemical spaces. We further visualize the molecule generation process in Fig. 3(right) with more examples deferred to Sec. F.5. It can be observed that our model explicitly preserves higher-order structures during the generation process.
@@ 4.2 Generic Graph Generation
To display the topology distribution learning ability, we assess HOG-Diff over four common generic graph datasets: (1) Community-small, containing 100 random community graphs; (2) Ego-small, comprising 200 small ego graphs derived from the Citeseer network dataset; (3) Enzymes, featuring 587 protein graphs representing tertiary structures of enzymes from the BRENDA database; and (4) SBM, a larger-scale stochastic block model dataset. The SBM benchmark is presented separately in Sec. F.2 due to its distinct evaluation protocol. Intermediate higher-order skeletons are obtained via 3-simplicial complex filtering. We employ the same train/test split as Jo et al. (2022) for a fair comparison. Maximum mean discrepancy (MMD) is used to quantify the distribution differences across key graph statistics, including degree (Deg.), clustering coefficient (Clus.), and 4-node orbit counts (Orbit). A low MMD signifies a close alignment between the generated and evaluation datasets, suggesting superior generative performance. We also report the average MMD across all metrics as an overall indicator.
We compare the following graph generative models: GraphRNN (You et al., 2018) is an autoregressive model, while GraphVAE (Simonovsky and Komodakis, 2018), GNF (Liu et al., 2019), and GPrinFlowNet (Mo et al., 2024) are one-shot models. GraphAF, GraphDF, EDP-GNN, SPECTRE, GDSS, and DiGress were introduced previously. The results in Tab. 2 show that HOG-Diff is not only suitable for molecular generation but also proficient in generic graph generation, demonstrating its ability to capture the intricate topological interdependencies effectively.
Table 3: Quantitative evaluation of higher-order topology preservation using Curvature Filtrations. The best results are in bold.
Dataset Method $\kappa_{FR}$ $\kappa_{OR}$
QM9 GDSS 0.925 0.601
DiGress 0.251 0.343
DeFoG 0.177 0.286
Cometh 0.216 0.314
HOG-Diff 0.077 0.098
ZINC 250k MiCaM 9.436 7.251
GDSS 1.781 1.331
DeFoG 0.728 0.498
HOG-Diff 0.190 0.098
MOSES DiGress 0.260 0.223
HOG-Diff 0.159 0.183
GuacaMol DiGress 0.862 0.075
HOG-Diff 0.083 0.057
SBM DeFoG 3.998 5.206
Cometh 2.748 3.938
HOG-Diff 1.453 3.740
Comm. -small GDSS 12.515 14.949
HOG-Diff 6.734 6.522
Enzymes GDSS 13.031 12.399
HOG-Diff 12.114 9.830
Ego -small GDSS 2.311 1.301
HOG-Diff 1.679 1.042
4.3 Topological Preservation Analysis
To further verify the capability of HOG-Diff in preserving higher-order structures, we conduct a quantitative evaluation using Curvature Filtrations (Southern et al., 2023). Unlike simple graph statistics, curvature filtrations combine discrete curvature notions with topological data analysis (TDA) to capture multi-scale topological features.
Specifically, we measure the distance between the mean persistence landscapes of generated graphs and the test set, using two filtration functions: (1) Balanced Forman–Ricci curvature $\kappa_{FR}$, which focuses on edge-based local clustering and cycles, and (2) Ollivier–Ricci curvature $\kappa_{OR}$, which captures global geometry and transport properties via Wasserstein distance. A lower distance indicates a generated distribution that is topologically closer to the test set.
Fig. 1 (right) reports the $\kappa_{FR}$ results, while Tab. 3 provides the full evaluation for both $\kappa_{FR}$ and $\kappa_{OR}$ across all molecular and generic datasets. HOG-Diff consistently achieves the lowest distance scores across both molecular and generic graph datasets. Notably, on complex molecular datasets, our method outperforms strong baselines by a significant margin, suggesting that while baseline models may capture basic chemical validity, HOG-Diff is superior in reconstructing the intrinsic topological backbone and higher-order geometric relations. Even on generic graphs like SBM and Enzymes, which possess distinct community structures, HOG-Diff maintains better topological fidelity, verifying that the performance improvement stems from the successful preservation of higher-order structures.
4.4 Ablations: Topological Guide Analysis
@@
Throughout the experiments, we observe that HOG-Diff exhibits superior performance on complex molecular datasets, but comparatively modest results on the Ego-small dataset. Statistics and visualizations in Apps. E and F.5 reveal that Ego contains the fewest higher-order structures among the datasets analyzed, suggesting that the choice of guide plays a pivotal role in the effectiveness of generation. To validate this hypothesis, we conduct further ablations using different types of topological information as guides.
In the sampling procedure, we further evaluate peripheral structures (Periph.), obtained by removing cell components, as guides. As shown in Tab. 4, both peripheral and noise guides underperform cell-based guides, providing empirical support for the tighter reconstruction error bound established in Theorem 4. These results indicate that certain topologies, particularly cells, are more effective in guiding generation, highlighting the importance of selecting appropriate topological structures to steer toward meaningful outputs. Moreover, this finding suggests that guides could serve as diagnostic tools for assessing whether specific topologies are essential to a given architecture. Systematic analysis of different guides promises to enhance interpretability, clarify how topologies affect generation, and inform the design of more effective graph generative models.
@@
Noise 91.52 0.829 0.0015
QM9 Peripheral 97.58 0.305 0.0009
Cell 98.74 0.172 0.0003
ZINC 250k Noise 96.78 1.665 0.002
Peripheral 97.93 1.541 0.002
Cell 98.56 1.533 0.001
5 Related Work
@@
Graph Generative Models. Early work on graph generative models dates back to random network models (Barabási and Albert, 1999), which offer foundational insights but are too simplistic for capturing real-world graph distributions. Recent advances in generative models have leveraged the power of deep neural networks, significantly improving the ability to learn graph distributions. Notable approaches include autoregressive or simultaneous models based on techniques such as variational autoencoders (VAE) (Jin et al., 2018; Simonovsky and Komodakis, 2018), recurrent neural networks (RNN) (You et al., 2018), normalizing flows (Zang and Wang, 2020; Shi et al., 2020; Luo et al., 2021), and generative adversarial networks (GAN) (Martinkus et al., 2022).
@@
A breakthrough in graph generative models has been marked by the recent progress in diffusion-based generative models (Niu et al., 2020; Ho et al., 2020; Song et al., 2021). Recent models employ various strategies to enhance the generation of complex graphs, including capturing node-edge dependency (Jo et al., 2022), addressing discretization challenges (Vignac et al., 2023; Huang et al., 2023; Siraudin et al., 2025), exploiting low-to-high frequency generation curriculum (Mo et al., 2024; Huang and Birdal, 2025), and improving computational efficiency through low-rank diffusion processes (Luo et al., 2023a). CatFlow (Eijkelboom et al., 2024) and DeFoG (Qin et al., 2025) adopt flow matching as an alternative to diffusion, achieving more efficient generation. Recent studies have also enhanced diffusion-based generative models by incorporating diffusion bridge processes, i.e., processes conditioned on the endpoints (Wu et al., 2022; Jo et al., 2024; Boget et al., 2025). Despite these advances, existing methods either overlook or inadvertently disrupt higher-order structures during graph generation, or struggle to model topological properties, as denoising noisy samples does not explicitly preserve the intricate structural dependencies required to generate realistic graphs. This highlights the need for a graph-friendly diffusion framework that explicitly learns higher-order topology, preserves meaningful intermediate states and trajectories, and avoids inappropriate noise injection.
@@ 6 Conclusion
We introduce HOG-Diff, a coarse-to-fine generation framework that exploits higher-order topology as explicit guidance. By employing cell complex filtering and diffusion bridges, HOG-Diff decomposes graph generation into structured, easier-to-learn, topology-aware substeps and preserves meaningful intermediate states. Our theoretical analysis supports its advantages over classical diffusion approaches, including improved training dynamics and a tighter reconstruction error bound. Extensive experiments on eight molecular and generic graph benchmarks further validate these gains. Overall, this work represents an important step toward topological diffusion models, highlighting the importance of higher-order features that prior methods overlook and leaving ample room for future work.
Limitations and future work
The performance of HOG-Diff depends on the presence of explicit higher-order structures, which certain graph types might lack. Nevertheless, higher-order motifs are often deducible in realistic graphs. Our future work will explore various filtering mechanisms beyond CCF. We will adaptively determine the order (dimension of the higher-order skeleton).
Acknowledgments